A​ 95% confidence interval is equivalent to which of the following hypothesis​ tests?

Choose the correct answer below.

A.

A​ two-tailed test with a significance level of 0.05

B.

A​ two-tailed test with a significance level of 0.10

C.

A​ right-tailed test with a significance level of 0.05

D.

A​ left-tailed test with a significance level of 0.05
Answer #1

Solution:
Given in the question
95% confidence interval so Confidence level = 0.95
Level of significance = 1 - confidence level = 1 - 0.95 = 0.05
alpha/2 = 0.05/2 = 0.025
So A 95% confidence interval is equivalant to two-tailed test with a significance level of 0.05.
So its answer is A. i.e. two tailed test with a significance level of 0.05